# Time-tagging data acquisition system for testing superconducting electronics based on an RFSoC and custom analog frontend R. A. Foster,11footnotetext: Corresponding author.S. Kandeh O. Medeiros

### v1

@@ 5 Conclusion and outlook
Although JJs are more mature than superconducting nanowires, and can more readily take advantage of specialized high-speed CMOS testing equipment like bit-error-rate testers, the sparse data acquisition capabilities of this firmware and ability to simultaneously control and monitor many inputs and outputs could still be useful to diagnose failures of circuits at high clock rates at a low cost.
Second, the timetagging sparse data acquisition could be useful for detector readout.
When superconducting nanowire single photon detectors (SNSPDs) are used as particle detectors [10], the pulse shape of the detection waveform may vary depending on the type of particle absorbed [37], so saving the actual waveform captured by the detector can provide insight into the detector performance.
In addition, photon-number-resolution is also possible by measuring the pulse height of an SNSPD biased through an impedance-matched taper [38].
The high sensitivity of the analog frontend would enable the data acquisition system to discriminate between single- and multi-photon events for an impedance-matched SNSPD.
@@
Even though an RFSoC was used, virtually all of the firmware is written to be device-agnostic, and could be used in other FPGA s, with the only significant area requiring modification being the dataconverter (RFADC/RFDAC) interface.
In the future, we plan to use this equipment to study the error rate of more complex superconducting circuits over various operating conditions such as temperature and external magnetic field.
@@ Acknowledgments
The authors thank Camron Blackburn, Malick Sere, and Donnie Keathley for helpful comments and feedback during the preparation of this manuscript. The authors thank Dr. Stephen Nagle of MIT TJ Rogers Laboratory for his assistance and guidance with assembly of the analog frontend. This work was funded by the DOE Office of Science Research Program for Microelectronics Codesign through the project “Hybrid Cryogenic Detector Architectures for Sensing and Edge Computing enabled by new Fabrication Processes” (LAB 21-2491). RF acknowledges funding from the Alan McWhorter fellowship. OM acknowledges funding from NDSEG fellowship. AS acknowledges funding from the NSF GRFP.

### v4

@@ 5 Conclusion and outlook
This opens up the possibility of using it to test RSFQ and AQFP logic circuits.
Although JJs are more mature than superconducting nanowires, the ability to simultaneously control and monitor many inputs and outputs could be very useful, particularly in situations where specialized high-speed CMOS testing equipment like the bit-error-rate tester used in [4] make it difficult to fully evaluate the possible input space of a circuit.
In addition to electronic circuits, the timetagging acquisition aspect of this work could be useful for detector readout.
Indeed, a similar scheme is already used by [30].
However, the applicability of this work to frequency-multiplexed readout is fairly limited, since the firmware would need to be heavily modified to introduce polyphase filter banks to perform demodulation of frequency-multiplexed resonator signals [43], and one would need to adjust the parameter configuring the number of channels for the sample discriminator to match the number of resonators.
For this reason, existing specialized FPGA-based solutions for readout of frequency-multiplexed superconducting detectors would be more appropriate for these applications [27, 28, 44].
However, the approach of saving the raw signal from the detector using the trigger-sharing mechanism presented in this work may still be useful for detector analysis, as it allows for recording of a minimal amount of data for subsequent postprocessing that would be impractical to run in real time.
For SNSPDs used as particle detectors [10], this may provide insight into what caused the detector to fire in the first place, because the pulse shape of the detection waveform varies depending on the type of particle absorbed [45].
It could also be useful for mitigating time-walk effects at very high count rates in SNSPDs or other current-biased detectors or transducers [46].
In addition, the high sensitivity of the analog frontend could enable the data acquisition system to discriminate between single- and multi-photon events for an impedance-matched SNSPD [47].
@@
Even though a ZCU111 RFSoC was used, virtually all of the firmware is written to be device-agnostic, and could be used in other FPGA s, with the only significant aspect requiring modification being the dataconverter (RFADC/RFDAC) interface.
It should be noted that for a smaller number of channels, products like the RFSoC4x2 from RealDigital may be a more cost-effective option if one does not require as many RFDACs or RFDAC operation down to DC.
In the future, we plan to use the data acquisition system presented in this work to study the error rate of more complex superconducting circuits over various operating conditions such as temperature and external magnetic field.
@@ Acknowledgments
The authors thank Camron Blackburn, Malick Sere, and Phillip Donnie Keathley for helpful comments and feedback during the preparation of this manuscript. The authors thank Stephen Nagle of MIT TJ Rogers Laboratory for his assistance and guidance with assembly of the analog frontend. This work was funded by the DOE Office of Science Research Program for Microelectronics Codesign through the project “Hybrid Cryogenic Detector Architectures for Sensing and Edge Computing enabled by new Fabrication Processes” (LAB 21-2491). RF acknowledges funding from the Alan McWhorter fellowship. OM acknowledges funding from NDSEG fellowship. AS acknowledges funding from the NSF GRFP.